OpenStudy (anonymous):

Solve for x: 5^x * 2^(x-1) = 100

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

guess last answer didn't help want to try again?

OpenStudy (anonymous):

I tried but I wasn't able to cancel out the logs to get the x value

OpenStudy (anonymous):

there is no cancelling it is a raft of algebra

OpenStudy (anonymous):

Could you show me how it's done. The first part I understand

OpenStudy (anonymous):

did you understand how to get \[x\ln(5)+(x-1)\ln(2)=\ln(100)\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

k then remember that although \(\ln(x)\) is a function, \(\ln(2), \ln(5),\ln(100)\) are numbers i.e. constants

OpenStudy (anonymous):

so it will be like solving \[2a+b(x-1)=c\] for \(x\) i can walk you through it if you like, but it is algebra from here on in

OpenStudy (anonymous):

oops i meant it will be like solving \[ax+b(x-1)=c\] for \(x\)

OpenStudy (anonymous):

Could u please go through it

OpenStudy (anonymous):

\[ax+b(x-1)=c\\ ax+bx-b=c\\ ax+bx=c+b\\ (a+b)x=c+b\\ x=\frac{c+b}{a+b}\]

OpenStudy (anonymous):

So we plug in the values to get x?

OpenStudy (anonymous):

repeat with \[\ln(2)x+\ln(5)(x-1)=\ln(100)\] it is identical

OpenStudy (anonymous):

as a matter of fact, now that we have it with \(a,b,c\) we can go right to \[x=\frac{\ln(100)+\ln(5)}{\ln(2)+\ln(5)}\]

OpenStudy (anonymous):

if you want a decimal, use a calculator

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

also if you like you could use properties of the log to rewrite as \[x=\frac{\ln(500)}{\ln(10)}\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Sorry to bother you but I've you plug in the x value (2.6989) in the original equation it doesn't equal 100

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

I found the error it should be In 200 instead of In 500

OpenStudy (anonymous):

Anyways thanks for your help. Now I understand how its done